OpenStudy (firejay5):

I don't know how to show my work on this and need help on it. Medal will be rewarded for work accurate, correct, and explained properly.

OpenStudy (anonymous):

do you mean on this site?

OpenStudy (firejay5):

Question and Answers

OpenStudy (anonymous):

well tell me how you got to those answers

OpenStudy (firejay5):

I took a stab at it, because our classes don't know how to do them well

OpenStudy (anonymous):

Well...unfortunately angles were my worst math subject....ever. I also haven't used them in years so although I could take a shot at it for you it would probably be better if you let someone else come along. If nobody else does message me and I'll try my best to explain it to you.

OpenStudy (firejay5):

I don't get it @jim_thompson5910 I mean to show my work

jimthompson5910 (jim_thompson5910):

angles 3, 4 and 5 are supplementary, which means they add to 180 degrees

jimthompson5910 (jim_thompson5910):

so... (angle 3) + (angle 4) + (angle 5) = 180 (x+18) + (x+15) + (4x+3) = 180 x+18 + x+15 + 4x+3 = 180 I'll let you solve

OpenStudy (firejay5):

I understand that

jimthompson5910 (jim_thompson5910):

ok what do you get when you solve for x

OpenStudy (firejay5):

24

OpenStudy (firejay5):

Okay I figured it out @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good